### v3

# GenAudit: Fixing Factual Errors in Language Model Outputs with Evidence

Abstract
LLMs can generate factually incorrect statements even when provided access to reference documents. Such errors can be dangerous in high-stakes applications (e.g., document-grounded QA for healthcare or finance). We present GenAudit — a tool intended to assist fact-checking LLM responses for document-grounded tasks. GenAudit suggests edits to the LLM response by revising or removing claims that are not supported by the reference document, and also presents evidence from the reference for facts that do appear to have support. We train models to execute these tasks, and design an interactive interface to present suggested edits and evidence to users. Comprehensive evaluation by human raters shows that GenAudit can detect errors in 8 different LLM outputs when summarizing documents from diverse domains. User studies demonstrate that using GenAudit can substantially improve the performance of humans at finding errors in LLM-generated summaries. We release our tool (GenAudit) and fact-checking model for public use. [^1] https://genaudit.org
1 Introduction
Figure 1: An illustration of GenAudit’s user interface and sample predictions. Reference document (a clinical transcript) is on the left and the generated text to be fact-checked is on the right (generated by querying any LLM, but manually entered here for ease of illustration). Spans in the text which are not supported or are contradicted by the reference are highlighted in red, with suggested replacements in green. As the user moves to any line in the generated text, evidence found for all facts in it are highlighted using blue links. Evidence and error predictions shown here are made by a fine-tuned Flan-UL2 model backend.
LLMs can produce factually incorrect or unsubstantiated statements (Li et al., 2023; Min et al., 2023), even when they are explicitly provided relevant context such as documents (Adams et al., 2023; Sadat et al., 2023). Incidentally, such document-grounded generation is often involved in high-stakes usage scenarios where factual correctness is (especially) paramount. For example, a doctor using an LLM to summarize a patient’s medical history (Adams et al., 2023; Kanwal and Rizzo, 2022) might make an incorrect decision if the generated summary contains errors. Manually verifying LLM outputs in such settings is therefore prudent, but also time-consuming and so undercuts the motivation for using language technologies in the first place. This motivates the need for a system that can assist users in efficiently verifying LLM output.
To this end, we introduce GenAudit, a tool for fact-checking LLM responses in document-grounded tasks such as summarization and question answering. Given a document and an LLM-generated output conditioned on the same, GenAudit (i) locates factual errors in the output text and proposes edits to fix them, and (ii) displays evidence to support facts in the (potentially edited) text. The system consists of two components: an interactive interface which presents evidence and edit suggestions for the user to act upon, and a bespoke backend model (fine-tuned LLM) capable of producing edits and identifying evidence. The interface allows the user to make edits to the LLM-generated text, and then observe updated predictions from the fact-checking model. Notably, in addition to supporting the task of fact-checking itself, the interface can also be used as a tool to evaluate and compare different backend fact-checking models, collecting data on human edits to fine-tune better models, and carry out counterfactual testing of fact-checking models by editing source documents.
We designed and evaluated different models to generate the fact-checking predictions for the tool, including fine-tuned and few-shot prompted LLMs. We treat this as a sequence-to-sequence task: Given an input document and a claim sentence, the model is required to simultaneously generate the sentence ids in the document which provide evidence, and a revised version of the claim which fixes any factual errors. We used data from the USB benchmark (Krishna et al., 2023) to train and evaluate models on the fact-checking tasks. We found that fine-tuned open-source LLMs perform better than few-shot prompted ChatGPT and GPT-4 models, at least when evaluated on an in-domain held-out test set.
[…]
Human evaluations also showed that GenAudit can be used to verify summarization outputs in different domains, including clinical conversations, news articles and social media posts. This is despite the fact-checking model being trained only on Wikipedia data. GenAudit successfully identified errors in outputs from 8 different LLMs including Mistral-7B (Jiang et al., 2023), LLama2-70B (Touvron et al., 2023), Gemini-pro (Team et al., 2023) and GPT-4 (Achiam et al., 2023) in human evaluation. Observed precision ranged between $79-100\%$ while, recall ranged from $23-57\%$ for different generation models.
We conducted a user study to measure the impact of assistance from GenAudit on the performance of humans trying to find errors in LLM-generated summaries. Each user fact-checked a number of summaries of news articles against the corresponding references, and for half of them they were provided assistance by GenAudit. We found that the percentage of erroneous words detected by the users was $62.2\%$ when they did not use GenAudit, and it increased to $79.4\%$ when they used GenAudit. This demonstrates the ability of the tool to boost performance of humans at fact-checking.
Users fact-checking LLM outputs would not want to miss any factual errors, and may prefer that that most potential errors are highlighted for review, even if some of them are false alarms. In terms of modeling, this translates to giving a higher importance to getting better recall than precision at error-detection. We therefore introduce a decoding algorithm for fact-checking models which generate revised/fixed versions of claims, which can increase the recall of error detection while minimizing drop in precision. This approach entails intervening at time-steps where the output probabilities fall below a threshold $\tau$ to select alternate decoding paths. Varying $\tau$ allows us to suggest more or fewer edits according to user’s preference, effectively trading recall against precision. This approach produces a better precision-recall frontier compared to baselines.
Input:
You are provided a document and its summary. The summary may potentially contain factual errors. The last sentence of the summary is marked as a claim. Find all sentences in the document providing evidence for the claim, and then revise the claim to remove or replace unsupported facts.
DOCUMENT: SENT0 Micheal Ward SENT1 Early life. SENT2 Micheal Ward was born in Spanish Town, Jamaica on 18 November 1997. SENT3 His mother was 18 years old when he was born. SENT4 He has three sisters. … SENT17 Ward’s breakout year came in 2019, when he starred as Jamie in Netflix’s revival and third series of "Top Boy". SENT18 He also appeared in a leading role in the film "Blue Story" in the same year. SENT19 The film received critical acclaim, and Ward won the BAFTA Rising Star Award for his performance. …
SUMMARY: Micheal Ward (born 18 November 1997) is a Jamaican-British actor and former model.
CLAIM: His films include "Blue Story" (2018) and "The Old Guard" (2020).
Output:
EVIDENCE: SENT18
REVISION: His films include "Blue Story".
Table 1: Sample datapoint with input-target formatting from the USB dataset
Our contributions are as follows:
  • We present GenAudit, a tool to assist fact-checking LLM outputs in document-grounded tasks. The tool identifies and fixes errors, and highlights evidence for claims.
  • We create fine-tuned LLMs which serve as backend models for fact-checking and evaluate them comprehensively on the task of fact-checking summaries generated by 8 LLMs for documents from 3 domains.
  • We demonstrate via a user study, that using assistance from GenAudit improves performance of humans at finding errors in LLM-generated summaries.
  • We present and evaluate a custom decoding algorithm that allows one to improve error detection recall while incurring a smaller drop in precision than baselines.
2 Background
We fine-tune LLMs to perform evidence extraction and claim editing, and use them as backend for GenAudit. In this section we provide an overview of the training dataset and models we use to power the underlying evidence extraction and factual error correction tasks.
2.1 The USB dataset
The USB dataset (Krishna et al., 2023) is composed of Wikipedia articles, their summaries and (human) annotations on them. The summaries have two versions: (i) An initial version which may have content that is unsupported by the article or contradicted by it, and (ii) An edited version which annotators have created by making minimal edits to the initial version to remove errors. Additionally, each sentence in the edited summary is linked to a minimal set of article sentences that provide sufficient evidence for all facts that it contains.
We format the dataset in a sequence-to-sequence format to use it for fine-tuning LLMs (Table 1). The input to the model starts with the task instruction. It is followed by the reference document where each sentence is prefixed by a sentence ID (e.g. SENT1, SENT2…). It is then followed by the summary sentences upto the sentence to be fact-checked (called the claim). The sentences preceding the claim are included so that relevant context from it (e.g. coreferences) can be used for better understanding of the claim. Finally, the claim is appended to the input. The target output consists of the two parts. The first part contains a list of sentence ids from the document which provide evidence for the claim, and the second part consists of a revised version of the claim which removes its unsupported information and replaces incorrect facts.
We use a custom split of the USB dataset for training and evaluating our model. We shuffle and divide the entire dataset into train, validation and test splits of size $94\%$, $3\%$ and $3\%$ of the full dataset. This differs from the original USB splits in two ways. First, the training split is much larger at $94\%$ instead of the original $40\%$. Second, the training split consists of articles from all 6 domains in the benchmark, whereas originally 2 of the domains where reserved as challenging OOD examples to occur in only the test set. The motivation for both of these changes is that we want to create a tool which generalizes to other diverse data beyond simply Wikipedia articles, and hence we train on as much and as diverse data as possible.
2.2 Reducing memory requirement for training
We aim to fine-tune large models for the fact-checking task, since they are more likely to generalize better to unseen domains due to internal knowledge. Additionally, we need to feed in the entire reference document to the model, which can be thousands of tokens long. Both these factors increase the memory requirement for training the models, and to address this challenge, we use low-rank adapters with 4-bit quantization (Dettmers et al., 2023). Low rank adapters (Hu et al., 2021) reduce the memory requirement during training by reducing the number of trainable parameters, thus reducing gradient computation. To reduce the sequence length in cases where the reference document (Wikipedia article) is too long, we iteratively drop sections in it which are not relevant to any sentence in the summary (i.e. do not provide any evidence for it). We follow this process until the input sequence length is reduced to within a maximum limit, which is kept the same regardless of what model we are fine-tuning. We also use gradient accumulation and gradient checkpointing (Chen et al., 2016) to reduce memory footprint for training. For more details, please see the Appendix.
Figure 2: Overview of the Genaudit interface. The reference document (A) is displayed alongside the model generated output (B), allowing for easy comparison. Users can initiate the fact-checking process by clicking the fact-checking icon next to each output sentence (C), which highlights potential factual inconsistencies. Evidence markers (D) indicate supporting or contradicting evidence from the reference document. Hovering over an evidence marker highlights the relevant section in the reference (E), and clicking on it scrolls to the corresponding location in the document. Hallucinated spans in the generated output are marked in red (F), with the system providing suggestions for real-time correction
3 System Overview
The GenAudit interface presented in Figure 2, allows users to interact with both the generated output and the corresponding reference document. This interface enables users to perform the following functions aimed at enhancing the factual accuracy of generated text.
  • Review generated output: The generated text (labeled B) is displayed adjacent to the reference document (labeled A). This side-by-side layout enables an intuitive comparison between the generated text and the document it is derived from, serving as a first step in the fact-checking process.
  • Fact-Check outputs: A key feature of the interface is the ability to fact-check generated outputs with a single action. Users can fact-check any sentence in the generated text by clicking the corresponding icon next to it (labeled C), which triggers an automated verification process that cross-references information in the generated output against the reference document. Users can also click on the “Fact Check All” button to trigger fact-checking of all generated sentences.
  • Highlight Evidence: The system displays evidence markers (labeled D) next to each generated sentence. These markers link to the corresponding sentences of the reference document that provide supporting or contradicting evidence. When users hover over these evidence blobs, the relevant sentence of the reference document is highlighted (labeled E) to make the connection clear. Clicking on the evidence marker automatically scrolls the reference text to bring the supporting content into view, helping users quickly locate relevant evidence.
  • Editable Text: In cases where factual inaccuracies are detected, GenAudit offers automatic edit suggestions to correct the hallucinated spans. Users can interact with these proposed changes by reviewing the marked hallucinated spans (highlighted in red and labeled F). The system allows users to accept or reject the suggested edits, enabling real-time refinement of the generated text.
4 Experiments
Model Error Identification Evidence
Recall Precision F1 Recall Precision F1
Finetuned decoder-only LLMs
Falcon-7B 69.03 61.54 65.07 59.85 54.23 56.90
Llama2-7B 74.85 39.19 51.44 68.03 68.47 68.25
Mistral-7B 80.53 73.34 76.77 72.25 86.66 78.80
Fine-tuned encoder-decoder LLMs
Flan-T5-XL 73.01 87.07 79.42 78.90 85.69 82.16
Flan-T5-XXL 80.38 84.50 82.39 81.46 85.96 83.65
Flan-UL2 76.47 87.44 81.59 80.56 86.42 83.39
Few-shot prompted proprietary LLMs
GPT-3.5-turbo (8shot) 38.79 48.57 43.13 51.79 45.15 48.24
GPT-4 (4shot) 37.98 63.89 47.64 74.42 38.52 50.76
Table 2: Performance of different models on the test split of the USB dataset for the tasks of (i) identifying erroneous words, and (ii) highlighting relevant evidence
Error Identification Replacements Evidence Extraction
BaseRate Recall Precision F1 AcceptedRecall Precision F1 Sufficient%
Aggregate 3.97 40.37 95.04 56.66 78.18 90.83 95.22 92.97 85.98
Summary generation models
Llama2-7B 4.29 30.21 89.29 45.15 66.67 90.71 96.12 93.33 86.65
Mistral-7B 1.99 23.83 92.00 37.86 40.00 91.43 95.80 93.57 87.59
Falcon-7B 21.84 47.95 97.46 64.28 86.36 84.65 87.08 85.85 76.77
Llama2-70B 3.29 43.38 95.93 59.75 100.00 91.98 92.09 92.04 86.10
Flan-UL2 9.68 34.04 96.04 50.26 80.00 90.18 93.96 92.03 84.16
Gemini-pro 1.80 27.75 78.69 41.03 66.67 91.27 96.98 94.04 86.68
GPT-3.5-turbo 1.10 29.06 89.47 43.87 75.00 92.32 97.23 94.71 88.09
GPT4 2.53 56.77 100.00 72.42 87.50 90.39 97.07 93.61 85.49
Datasets
XSum 4.54 55.00 98.69 70.64 77.78 94.81 97.92 96.34 92.83
ACIBench 2.73 44.04 90.65 59.28 88.89 87.57 93.30 90.34 80.84
Reddit 4.88 22.52 92.41 36.22 60.00 91.93 95.50 93.68 86.55
Table 3: Results from human evaluation of GenAudit predictions (using fine-tuned Flan-UL2 model). Aggregate results show performance on all evaluated document-summary pairs, whereas other rows show performance (a) on summaries generated using individual LLMs, or (b) on documents from specific datasets. BaseRate represents the percentage of tokens in summaries that are hallucinations. Sufficientrepresents the percentage of summary sentences for which the evidence shown by the model was sufficient for deciding factuality.
We use the USB dataset for training/prompting and evaluating 8 different models for two factchecking tasks. For the evidence extraction task, we report precision, recall and F1 score as in a binary classification task where given a claim and reference document, each sentence in the reference is classified as relevant evidence or not. To evaluate the model’s ability to remove errors, we compare the words removed/replaced by the model vs those removed in the ground truth revision. Given the original claim and a revision, we tokenize each into words and compute the diff between them (using Python’s showdiff library). The words in the claim that are removed/replaced in the revision are tagged as incorrect and the remaining words are tagged as correct. We use the ground truth revision in the dataset to compute the ground truth tags, the model-generated revision to compute the predicted tags, and compute the corresponding precision, recall and F1 scores. Notably, it is difficult to compare the replacement text proposed by the model with ground truth replacements automatically, since the underlying text span being replaced must match exactly to make an aligned comparison. This requires human evaluation which we discuss in the next section.
We fine-tune and evaluate 6 different models. These include decoder-only models from the Falcon (Almazrouei et al., 2023), Llama2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) series, and encoder-decoder Flan-T5 models  (Chung et al., 2022; Tay et al., 2022). Finally, we also use OpenAI’s GPT-3.5-turbo and GPT-4 models for the task via few-shot prompting, using 8 and 4 exemplars respectively.
We find that there is large variation in performance of decoder-only LLMs (Table 2). Llama2 outperforms Falcon in evidence extraction, but underperforms it in the claim editing task, due to its low precision. Mistral outperforms both Falcon and Llama2 models by a large margin. There is relatively less variation in performance of the three encoder-decoder models. Flan-T5-XXL and Flan-UL2 models perform the best, with the former providing better recall and the latter providing better precision on both tasks. Few-shot prompted GPT models perform worse than all fine-tuned models on both error identification and evidence extraction.
5 Human Evaluation
In the previous section we saw that models fine-tuned on the USB dataset perform well when evaluated on its test split. However, this does not imply that they would also perform well when deployed in diverse out-of-domain scenarios. Two types of domain shift can occur here. The first is a change in the domain of reference documents used for fact-checking. USB consisted of Wikipedia articles only, but we would ideally want a finetuned model to work with other document types such as news articles or meeting transcripts. Another sort of domain shift is the specific model generating the content to be fact-checked. USB consists only of claims written by humans, but we would want models to detect and fix errors in content generated by a arbitrary LLMs.
We run experiments to evaluate the performance of GenAudit when fact-checking summaries generated by different models for documents sampled from different domains. We include a diverse set of open-source models, including three decoder-only 7B parameter LLMs (Mistral (Jiang et al., 2023), Llama2 (Touvron et al., 2023), Falcon (Almazrouei et al., 2023)), one large 70B parameter model (Llama2), and one encoder-decoder model (Flan-UL2 (Tay et al., 2022)). As proprietary API-based models, we use GPT-3.5-turbo, GPT-4, and Gemini-pro models for summary generation. We then use the Flan-UL2 model finetuned on USB to fact-check the generated summaries.
We select documents for summary generation from the following three datasets.
XSum (Narayan et al., 2018)  A summarization dataset consisting of BBC news articles covering diverse topics and events.
ACI-Bench (Yim et al., 2023)  A dataset for summarization of patient visits to the doctor comprising transcripts of doctor-patient encounters.
Reddit-TIFU (Kim et al., 2019)  A dataset consisting of posts from the online discussion forum Reddit in which users narrate personal day-to-day experiences.
We randomly select 30 documents from each of the three datasets for which to generate summaries. For the Reddit-TIFU dataset, we manually filtered out examples containing profanity or sexually explicit content. While generating summaries with open-source models, we decode using top-$p$ nucleus sampling (Holtzman et al., 2019) from the output token distribution with a top-$p$ value of $0.9$ and a temperature of $1.0$.
We hired annotators via Upwork, [^3] https://www.upwork.com/ and instructed them to evaluate all edits suggested by the fact-checker, accept those that fix legitimate factual errors, and mark incorrect suggestions. Annotators were also instructed to find any missing errors in summaries which were not highlighted by the system, and to fix them by making minimal edits.
To provide feedback on the highlighted evidence, annotators provided binary (relevant/not relevant) feedback for each suggested evidence sentence in the summary. They were also instructed to consider if the highlighted evidence for each summary sentence is sufficient or not; of not, then they should mark additional source sentences which contain the missing evidence. Additionally, we asked annotators to flag incomprehensible summaries, which were then excluded from the analysis. For example, instead of a summary, sometimes Falcon-7B model outputs a continuation of instructions, such as “when creating a summary, use the information given and avoid superfluous details.” The Appendix includes additional evaluation details.
Results are shown in Table 3. We use the metrics described in Section 4 for rating suggested errors and evidence generated by GenAudit. On an aggregate level—across all domains and summary generation models—GenAudit identifies erroneous words with high precision ($95.04\%$) and moderate recall ($40.37\%$). Note that achieving high recall is challenging here given the low prevalence of erroneous words ($3.97\%$). With respect to evidence extraction, we observe high precision and recall ($95.22\%$ and $90.83\%$, respectively), suggesting that most evidence sentences highlighted by the model are useful for fact-checking the given claim, and only few evidence sentences are missed (not highlighted) by the model. We have provided some samples of model-predicted errors in the Appendix, with the full collection available online [^4] https://aftersunrise.pythonanywhere.com (anonymized) .
[…]
Among the datasets considered, GenAudit’s performance at error identification was best for XSum (news articles), followed by ACIBench (clinical conversations), and finally Reddit/TiFU (social media posts). While the precision stays above $90\%$ on all datasets, the recall ranges from $22.52\%$ on the Reddit dataset to $55.00\%$ on XSum. On the evidence extraction task, GenAudit achieves F1 scores of $90\%+$ on all three datasets.
[…]
Using the annotations collected above, we evaluate additional models on the error detection task. Few-shot prompted GPT-4 achieves better recall than Flan-UL2, while other fine-tuned models achieve lower recall (Table 4). All models achieve lower precision than Flan-UL2. Edits suggested by GPT-4 add about $8.8\%$ more words to them on average, while other fine-tuned models add a negligible amount, reflecting the tendency of GPT-4 to make substantial changes to text. Finally, we also evaluate the FAVA model trained by Mishra et al. (2024) for factual error correction. We see that the model achieves the lowest recall compared to all the models evaluated, with a slightly higher precision than GPT-4.
Model Recall Precision %Del %Add
Flan-UL2 40.37 95.04 1.69 0.18
Flan-T5-XL 25.75 74.23 1.38 0.12
Mistral-7B 35.08 45.24 3.08 0.16
GPT-4 (4-shot) 40.68 28.50 5.67 8.80
FAVA (Mishra et al., 2024) 14.18 31.34 1.80 0.43
Table 4: Performance of models fine-tuned by us on the USB dataset, few-shot prompted GPT-4, and the FAVA model (Mishra et al., 2024) at identifying erroneous words in model-generated summaries, along with the percentage of summary words deleted and added by their edits.
6 Impact on human performance at fact-checking
In the previous section, we asked humans to evaluate the correctness of the factual errors detected by GenAudit, as well as the relevance of evidence highlighted by it. While those experiments show how accurate GenAudit is, it does not reveal how helpful it would be to a human carrying out the fact-checking exercise by themselves. In this section, we measure how the fact-checking performance of a human user would change if they were given access to GenAudit.
We recruited human subjects to carry out fact-checking of summaries, and measured their performance with and without access to GenAudit predictions. In each task, users were given 2 article-summary pairs and were asked to find as many errors as they can in them. For one summary, they were shown the highlighted evidence and suggested edits from GenAudit, and for the other summary they weren’t shown the outputs of GenAudit. The rest of the interface was exactly the same. The time available to the human subject for fact-checking a summary can heavily influence their performance, and performance can perhaps increase indefinitely as more time is provided. To control for this aspect, we asked the subjects to carry out the fact-checking exercise under time constraints. We gave them 3 minutes to fact-check each summary.
The subjects were recruited using a qualification task where each candidate was given two article-summary pairs to fact-check — one with help from GenAudit and one without. The annotation exercise was attempted by $23$ candidates. We chose the top $4$ performing candidates, where their performance was measured by the F1-score of error detection.
To create the data for this experiment, we took articles from the XSUM dataset and generated summaries using GPT-3.5-turbo model. Since the rate of errors in summaries generated are quite low (ref. Table 3), most of the summaries won’t have any errors. This can be problematic, because we would like to have a relatively large number of errors to get a reliable estimate of subjects’ performance at error detection. Hence, we chose to inject errors in each summary synthetically by prompting GPT-3.5-Turbo again, following the process outlined in Laban et al. (2023) (more details are provided in the Appendix). This provided us with a larger pool of errors, as well as ground truth labels marking the erroneous spans in the summaries post-error-addition.
[…]
7 Improving Recall of Error Detection
Users fact-checking LLM outputs using GenAudit may give more importance to a higher recall than precision to be confident that most errors are highlighted for review, even at the cost of false positives. While it is always possible to increase recall by indiscriminately flagging additional text spans as errors, a naive strategy would lead to a large drop in precision. We propose a decoding algorithm for the fact-checking model which uses the output token probabilities to achieve a better precision-recall trade-off.
Our proposed approach (Algorithm 1) for increasing error detection recall relies on observing the probabilities of tokens generated as the revision by the fact-checker, and intervening at timesteps with low model confidence. Given a document $D$, claim $C$, and an initially generated revision $R=r_{1}r_{2}..r_{m}$, we find the first position $t$ where the probability assigned to token $r_{t}$ falls below a threshold $\tau$. At that timestep we then generate the token with the highest probability excluding $r_{t}$. We generate the remaining tokens (from $t+1$) as usual via greedy decoding to compute an alternate revision $R^{\prime}$. Given $R$ and $R^{\prime}$, assume the span $r_{k}r_{k+1}..r_{k+w}$ was replaced by $x_{1}..x_{q}$. We make the replacement in $R$ yielding $r_{1}...r_{k-1}x_{1}..x_{q}r_{k+w+1}...r_{m}$. We then repeat the process of finding low probability tokens and making edits for the remaining tokens. After each iteration of the while loop, the value of $(|R|-t)$ decreases by at least 1, which guarantees termination of the program.
Algorithm 1 Thresholded Edit
  Input: document $D$, claim $C$, predicted evidence $E$, predicted revision $R$, model $\mathcal{M}$, threshold $\tau$
  $Q=(D,C,E)$
  $t=1$
  while $t\leq|R|$ do
     //  run this loop until the timestep counter reaches end of string R
     $p_{1}p_{2}...p_{|V|}$ = ${\text{NextTokProb}}_{\mathcal{M}}(r_{1}..r_{t-1}\mid Q)$run M to get probability of tokens that can be generated at timestep t}
     if $p_{r_{t}}\leq\tau$ then
        //  if the probability of the token currently at timestep t in R is lower than tau, we replace it
        $r^{\prime}=\arg\max_{k}(p_{k}\mid k\neq r_{t})$replace it with the token with the next highest probability (call it r’)}
        $\text{prefix}=r_{1}r_{2}..r_{t-1}r^{\prime}$
        compl = ${\text{Generate}}_{\mathcal{M}}(\text{prefix}\mid Q)$use M to generate the full completion of the revision if we use r’ at timestep t}
        $R^{\prime}=\text{prefix}+\text{compl}$we call the new revision R’}
        $N_{del},N_{add},\text{repl}$ = $\text{{DiffAtPos}}_{t}(R,R^{\prime})$extract span-level change between R and R’ at timestep t (deletions/replacements)}
        $R=\text{prefix}+\text{repl}+r_{(t+N_{del})}...r_{|R|}$commit that change at t in R and discard any other changes observed later}
        $t=t+N_{add}$
     end if
     $t=t+1$advance the pointer and repeat the loop to find more spans to edit in R if they exist}
  end while
  Output: updated revision $R$
Increasing the value of $\tau$ in Algorithm 1 would lead to more edits being made to the claim, and vice-versa. We run the Flan-UL2 fact-checking model with different values of $\tau$ ranging from $0.0$ to $0.99$ and plot the resulting recall and precision at detecting errors annotated in the human evaluation experiment in Section 5. We find that using Algorithm 1, we are able to increase the recall from about $40\%$ to $60\%$, with a drop in precision from $95\%$ to $40\%$. Although there is a drop in precision, the drop is much lower than what one would get by using a simple randomized baseline. The baseline strategy we compare against is to boost recall by randomly selecting additional words (beyond the ones already predicted as erroneous by the model) and mark them as erroneous too. We compute the number of words that need to be selected to boost expected recall to a certain level, and the resulting drop in expected precision that it entails (see Appendix for derivation). The thresholding approach maintains a much higher precision with increasing recall compared to the baseline strategy, where the precision already falls to around $28\%$ when the recall increases to merely $43\%$ (Figure 3).
We also compare using the custom decoding strategy in Algorithm 1 with simply flagging additional tokens as non-factual if their probability of generation (by the fact-checking model) falls below a variable threshold. We see that this strategy performs worse than using Algorithm 1 (Figure 3). This suggests that post-hoc usage of token probabilities from the fact-checking model does not isolate the non-factual spans as well as active intervention during the decoding process as done in Algorithm 1.
Figure 3: Variation in precision and recall of error identification by a fine-tuned Flan-UL2 model when using thresholded editing (Algorithm 1), versus editing out additional tokens either at random or by selecting the ones with low probability.
8 Related Work
With the rapid rise in capabilities of large language models (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023), they have been widely used for assisting humans in content generation workflows such as writing summaries (Jung et al., 2023), or scientific articles (Singh et al., 2024). However, it has been known for a long time that such language models produce factually inaccurate statements, popularly known as hallucinations (Arthur et al., 2016; See et al., 2017; Cao et al., 2018). This leads to a lack of trust in the AI-generated assistance, and multiple works have noted this limitation (Qian and Wexler, 2024; Swaroop et al., 2024). To ensure factual accuracy of text, some prior works have designed intelligent user interfaces. Rony et al. (2020) introduced a tool which helps journalists fact-check claims made in debates by retrieving relevant evidence via web-search. In contrast, our work also predicts whether claims are true or false in addition to highlighting relevant evidence from reference document. Concurrent work from Fu et al. (2024) proposed a tool where claims in text are verified against structured data tables. In contrast, our system works with free-form text as reference, and does not need access to data tables whose creation can be complex and may capture only specific categories of information.
Multiple works in the natural language processing literature have introduced approaches to predict whether a generated summary contains any factual error, either using trained models (Kryściński et al., 2020; Goyal and Durrett, 2021), question-answering based approaches (Fabbri et al., 2022), or prompting LLMs (Laban et al., 2023). While these works predicted factual correctness with respect to a given source document, recent works have implemented fact-checking against a large corpus (Wikipedia) by combining evidence retrieval and factuality prediction (Kamoi et al., 2023; Min et al., 2023). Our effort goes beyond binary prediction of factual correctness, by also localizing the errors in the claims and fixing them via minimal editing. Mishra et al. (2024) also attempts to fix factual errors via editing, and we compared the performance of their released model with ours in Section 5.
Liu et al. (2023) and Krishna et al. (2023) introduced the DeFacto and USB datasets respectively with human annotations to train and evaluate models for revising incorrect claims, and extracting evidence from a reference document. While both datasets can potentially be used to train GenAudit backend models, we used the USB dataset because (1) it contains comprehensive evidence labels for all facts in the claim, and (2) it contains multi-sentence summaries, which are more common in practice. We extend these lines of work by contributing an interactive tool for fact-checking, and a comprehensive evaluation of the models trained on such data at fixing errors in modern LLM outputs with evidence.
9 Conclusion
We introduced GenAudit, a tool to assist users in fact-checking LLM generated outputs against inputs by presenting supporting evidence and highlighting (and fixing) errors. We trained models for fact-checking tasks which rival few-shot prompting of SOTA LLMs, and designed a web-interface for users to interact with. We evaluated GenAudit for fact-checking summaries generated by 8 LLMs for documents in 3 domains. We demonstrated, via a user study, that using GenAudit improves performance of humans at fact-checking summaries. Finally, we proposed a decoding algorithm for our fact-checking model to improve the recall of error identification while minimizing the cost in precision.
Acknowledgements
This work was supported in part by the National Science Foundation (NSF), grant RI: 2211954 and 2211955.
Appendix A Appendix
Appendix B Binary Classification of Factuality
In the previous sections, we evaluated the performance of GenAudit at localizing factual errors within text and suggesting edits. However, it can also be repurposed as a binary classifier which simply predicts whether a long-form generated text is factually consistent or not with respect to given reference. To do that, we simply declare a given passage of text as factually inconsistent with respect to reference document if GenAudit suggests any edit to any sentence in it.
We evaluate the performance of GenAudit on the SummEdits benchmark (Laban et al., 2023) which consists of document-summary pairs where the summaries potentially contain factual errors. The source documents are taken from 10 different datasets representing a diverse group including legal documents, scientific papers, emails etc. We tokenize the source document into individual sentences before passing it through the fact-checking model. GenAudit achieves a balanced accuracy score of $74.7$, outperforming many LLMs and traditional fact-checking methods, with the exception of Gemini-pro and GPT-4 (Table 5) [^5] Values taken from the official Github repository https://github.com/salesforce/factualNLG.
Model Balanced Accuracy
Human Performance 90.92
GPT4 82.06
Gemini-pro 75.49
GenAudit 74.75
Claudev21 74.36
Claudev2 73.58
ChatGPT 71.18
PaLM-bison 69.04
QAFactEval (Fabbri et al., 2022) 65.46
Llama2-13b 58.35
Mistral-7b 57.78
SummaCConv (Laban et al., 2022) 57.14
DAE (Goyal and Durrett, 2021) 55.17
Llama2-7b 50.36
Table 5: Performance of models on the SummEdits benchmark for binary classification of factuality. Here GenAudit uses the fine-tuned Flan-UL2 backend, whereas other LLMs are zero-shot prompted.
B.1 Details on human evaluation (for Section 5)
We engaged annotators via Upwork [^6] https://www.upwork.com/, leveraging their expertise to evaluate model-generated summaries against source documents. Candidates were chosen through a qualifying round, focusing on their ability to identify inaccuracies in summaries. Ultimately, two proficient proofreaders and fact-checkers were selected based on their performance on the qualifying task. Each annotator was tasked with annotating all summaries for half of the documents in each of the 3 datasets used (see Section 5). Annotators received compensation at an average rate of \$25 USD per hour for their contributions.
We use a slightly modified version of the GenAudit UI to collect the annotations, with two notable changes (Figure 5). First, we add a buttons next to every source sentence to provide a accept/reject feedback if the source sentence is highlighted as evidence for a summary sentence. For source sentences which are not highlighted as evidence, we provide an accept button to mark it as additional evidence if needed. Second, the UI enables cycling through multiple model-generated summaries for the same source document at once. This is done to save annotators’ time, since otherwise the annotators would have to read the source document again each time they get a summary for it in the sequence of annotation jobs. The annotators were instructed to mark the following categories of generated summary sentences as invalid which we excluded from our analysis: (i) truncated sentences, which occur due to the maximum decode length limit being reached. (ii) incomprehensible sentences, such as when the models generate instructions instead of summary (e.g. Falcon-7B once generated “when creating a summary, use the information given and avoid superfluous details.”)
B.2 Annotator Instructions (for Section 5)
The following protocol was provided to annotators for assessing both the supporting evidence for summary sentences, and their factual consistency:
Evidence Annotation:
a) Evaluate each summary sentence by reviewing all linked evidence from the source, marking them as either accepted (by clicking a tick) or rejected (by clicking a cross) based on their validity.
b) Examine the summary for unsupported information with respect to suggested evidence, If any information supporting the summary is found in the source but is not already supported by the suggested evidence, mark it as "new" evidence. Note that "new" evidence pertains only to instances where the summary includes seemingly unsupported information compared to the suggested evidence.
Summary Annotation:
a) First, accept all suggestions made by the tool that are valid. This encompasses deletions made by the tool of unsupported information and any recommended edits or replacements.
b) If there is incorrect or contradictory information in the summary not satisfactorily addressed by the tool then make minimal edits to the summary to align it with the source. If the edit is based on source information not already included in suggested evidence, label it as "new evidence."
c) For cases where the summary introduces unsupported information not addressed by the tool, remove corresponding segments of the summary without altering the evidence.
Note if a summary sentence is incomplete or incomprehensible, mark the sentence as INVALID.
B.3 Inter-annotator agreement (for Section 5)
For one-third of the documents in each dataset, we got their model-generated summaries annotated by both annotators to estimate inter-annotator agreement. The doubly-annotated data included $232$ summaries consisting of a total of $989$ sentences. We compute the agreement on error identification by comparing the words removed/replaced from the initial summary by the two annotators. The value of Cohen’s Kappa for this is $62.7$ indicating substantial agreement. Similarly, we compare the ratings (useful vs not) provided to each suggested evidence sentence by the two annotators, which yields a Cohen’s Kappa value of $58.56$ indicating moderate agreement.
B.4 Creating data for human fact-checking performance evaluation (for Section 6)
We sampled news articles randomly from the XSum dataset to use as reference text, and used GPT-3.5-Turbo to generate their summaries. We further used GPT-3.5-Turbo to modify the summaries generated in this first round by adding factual errors. We used the prompt template used by Laban et al. (2023) to generate the modified summaries with added errors. [^7] https://github.com/salesforce/factualNLG/blob/master/prompts/summedits/step2_inconsistent.txt The prompt template is shown in Figure 4.
Document:
[DOCUMENT]
Consistent Summary:
[CONSISTENT_SUMMARY]
Given the document and consistent summary above, generate one slightly modified
version of the summary such that the modifications introduce a factual
inconsistency. For example, you can modify a number, date, or entity, and negate
or modify a statement. Here are some rules to follow:
- The modifications should change at most 5-6 words from the original summary,
 and keep the rest the same.
- The modifications should change  different parts of the original summary.
- Your modifications should be challenging to detect: modify minimally while
 still introducing a factual inconsistency.
- The factual inconsistency you introduce should be subtle. For example if you
 replace an entity, make sure you replace it with another entity from the document.
- The modified summary should start with "[FIRST_THREE_WORDS] [...]", and end
with "[LAST_THREE_WORDS]"
Figure 4: Prompt template used for introducing factual errors in summaries used for human fact-checking performance evaluation (Section 6). Taken from Laban et al. (2023)
B.5 Derivation of baseline precision-recall trade-off
Assume a binary classification problem with a dataset of $T$ datapoints, out of which $P$ have positive labels and $N$ have negative labels. Let’s assume a prediction model assigns positive labels to $P^{\prime}$ datapoints, and achieves a recall of $\alpha$ and precision of $\beta$. We want to boost the recall to a target of $\alpha^{\prime}$ by flipping the predictions from negative to positive for some datapoints. In this section we derive the precision-recall trade-off achieved if we select the labels to flip uniformly at random, which is the baseline used in Section 7.
If we flip the labels of $k$ datapoints from negative to positive, the expected number of them which would be true positives will be
$\Delta=\frac{P(1-\alpha)}{(T-P^{\prime})}k$
For an expected target recall of $\alpha^{\prime}$, we get
$\alpha^{\prime}=\frac{\alpha P+\Delta}{P}$
$\implies\alpha^{\prime}=\alpha+\frac{k(1-\alpha)}{(T-P^{\prime})}$
$\implies k=\frac{\alpha^{\prime}-\alpha}{1-\alpha}(T-P^{\prime})$
The expected true positives is $\alpha^{\prime}P$, which yields the new expected precision $\beta^{\prime}$ to be
$\beta^{\prime}=\frac{\alpha^{\prime}P}{(P^{\prime}+k)}$
$\implies\beta^{\prime}=\frac{\alpha^{\prime}P}{P^{\prime}+\frac{(\alpha^{\prime}-\alpha)}{(1-\alpha)}(T-P^{\prime})}$
Note that $P=\gamma T$, where $\gamma$ is the base rate for positive class, and the number of predicted positive labels $P^{\prime}=\frac{\alpha P}{\beta}$. Substituting these, we get our final form
$\beta^{\prime}=\frac{\alpha^{\prime}\gamma}{\frac{\alpha\gamma}{\beta}+\frac{(\alpha^{\prime}-\alpha)}{(1-\alpha)}(1-\frac{\alpha\gamma}{\beta})}$
We showed the performance of this baseline against our proposed Algorithm 1 in Figure 3.
Figure 5: Interface used for collecting feedback on suggested evidence and edits from GenAudit. Annotators can accept/reject each suggested evidence sentence, and can also mark additional sentences as evidence if needed. Suggested edits can be accepted/rejected by clicking on the button on the top-right of the highlighted span. if needed, users can also make freeform edits to fix more errors. Annotators cycle through the summaries generated by different models, whose names are anonymized and their order is shuffled.
B.6 Training details
Each model fine-tuned on the USB dataset, was trained for 10 epochs and the best checkpoint was selected based on validation performance at the end of each epoch. The effective batch size was kept at 128. The optimizer used was 8-bit AdamW with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, without weight decay, and with a learning rate of $55e-$ 5. The maximum input and output sequence lengths were set to $3050$ and $150$ tokens respectively. We used 4-bit quantization with low-rank adapters (Dettmers et al., 2023) with parameters $r=8$ and $\alpha=32$
Each training run was carried out on 2×2$\times$2 × Nvidia A6000 GPUs each of which has 49GB of memory. We used gradient accumulation and gradient checkpointing to reduce the peak memory requirements during training. The duration of each training run varied depending on the model used but fell in the range of 30-70 hours. The implementation was done using the Huggingface transformers (Wolf et al., 2019), Deepspeed (Rasley et al., 2020), Peft (Mangrulkar et al., 2022) and Bitsandbytes [^8] https://github.com/TimDettmers/bitsandbytes libraries.
Table 6: Links to models used in Section 4 and Section 5
Table 7: Links to datasets used in this work
Model Prompt
GPT-4 You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim. Make sure to provide evidence for all the supported facts in the claim. Then, revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. Only make changes to the text of the claim when necessary. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
GPT-3.5-turbo You are provided a document and its summary. The summary may potentially contain facts which contradict with the document or are not supported by any evidence in the document. The last sentence of the summary is marked as a claim. Find and list sufficient sentences in the document to provide evidence for the claim, and then revise the claim to remove or replace facts which are not supported by the document or are contradicted by it. When you add new information to the claim, it must be only to fix a contradictory fact in the claim, and not for changing the style of the text.
Others You are provided a document and its summary. The summary may potentially contain factual errors. The last sentence of the summary is marked as a claim. Find all sentences in the document providing evidence for the claim, and then revise the claim to remove or replace unsupported facts.
Table 8: Prompts used for fact-checking using GPT models in zero-shot setting, and using other models with fine-tuning (as described in Section 4)
Model Prompt
Gemini-pro, GPT-3.5-turbo, GPT-4 Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document. Generate the summary in free-form text without using bullet points.
Others Generate a summary for the following document in brief. When creating the summary, only use information that is present in the document.
Table 9: Prompts used for summary generation using different models for human evaluation experiment (Section 5)
Figure 6: GenAudit suggested edits for a GPT4-generated summary of a news article from the XSum dataset. The first sentence contains a statement attributed to the UK which was actually made by the president of the European Union
Figure 7: GenAudit suggested edits for a GPT4-generated summary of a news article from the XSum dataset. Interestingly, The reference document does not contain the date on which the said statement was made by SFA. Interestingly though, GPT-4 almost got it right by sheer memorization. The statement was released on 17 November 2016 whereas GPT-4 mentioned 10 November 2016. (Ref: https://www.bbc.com/sport/football/38021627,https://www.bbc.com/sport/football/38019477)
Figure 8: GenAudit suggested edits for a Geminipro-generated summary of a Reddit post where a person describes fire alarm going off at a workplace due to smoke from burnt toast. The summary suggests that there was a fire caused which doesn’t seem to be the case from the reference.
Figure 9: GenAudit suggested edits for a GPT4-generated summary of a conversation between a doctor and patient. Here, the doctor briefly mentions that the patient’s blood sugar problems may be caused by eating chocolates, but they don’t suggest that such unhealthy consumption has increased recently (as the summary claims).